4. Some students prepare 3 dfferent snacks. They make pound bags of nut mix, pound bags of cherries, and pound bags of dried fruit. If they buy 3 pounds of nut mix, 5 pounds of cherries 4 pounds of dried fruit, how many of each type of snack bag will they be able to make?
Answer #1

they made 1/8 pound bags of nut mix

and they bought 3 pounds of nut mix

so bags of nut mix = 3 / 1/8 = 24 bags

they made 1/4 pound bags of cherries

and they bought 5 pounds of cherries

so, bags of cherries = 5/1/4 = 20 bags

they made 1/6 pound bags of dried fruits  

and they bought 4 pounds of dried fruits

so, bags of dried fruits = 4 / 1/6 = 24 bags